OpenStudy (anonymous):

Verify that cos x + pi/2 = - sin x

OpenStudy (anonymous):

will give medal to answer

OpenStudy (anonymous):

cos (pi/6) +x = sin (pi/6), - (sqrt 3)/2 + (sqrt 3)/2 + x = sin (pi/6) - (sqrt 3)/2, x= sin (pi/6) - sin (pi / 3), x= -sin (pi/6). Therefore, cos x minus sin x equals sin x.